What relation exist between the range covered by two projectiles thrown up with the same initial velocities at angles of projection that are complementary
1
Expert's answer
2020-10-29T07:37:10-0400

Let 1 represent the first body and 2 represent the second body


"u_1 = u_2 = u"


"\theta_1 + \theta_2 = 90°" (complementary angles)

"\therefore \theta_2 = 90° - \theta_1"


For body 1,


"R_1 = \dfrac{{u_1}^2\, sin2\theta_1}{g}"


"R_1 = \dfrac{u^2\, sin2\theta_1}{g}"


For body 2;


"R_2 = \dfrac{{u_2}^2\, sin2\theta_2}{g}"


"R_2 = \dfrac{{u}^2\, sin2(90-\theta_1)}{g}"


"R_2 = \dfrac{{u}^2\, sin(180-2\theta_1)}{g}"


"sin(180-\theta) = sin \theta\\
\textsf{Then, }sin(180 - 2\theta) = sin2\theta"


"\therefore R_2 = \dfrac{{u}^2\, sin2\theta_1}{g}"


"R_1 = R_2"


"\therefore" The range covered by two projectiles thrown up with the same initial velocities at angles of projection that are complementary is equal